A particle moving in one dimension with a constant acceleration of 2m/s² is observed to cover a distance of 5 m during a particular interval of one second. The distance covered by the particle in the next one second interval is (in metres) :
a)5 , b)6, c)7(answer) , d)10 , e)12
How do you get this answer? Thanks!

Grade:11

1 Answers

Aditya
7 years ago
According to question particle covers 5m in 1 second so it`s speed Is 5mps. Applying v = at the instant at which speed is 5 is given by 5 = 2 * t So t = 5/2s . The next second means 5/2 + 1 = 7/2 s it`s speed will be v = 2*(7/2) = 7metres in 1 second